OpenStudy (mimi_x3):

At a point P on the parabola x^2=4ay, a normal PK is drawn. From the vertex O a perpendicular OM is drawn to meet the normal at M. Show that the equation of the locus of M as P varies on the parabola is x^4 -2ax^2y + x^2y^2 - ay^3 = 0

OpenStudy (anonymous):

\[x^2 = 4ay\]  Differentiating with respect to x \[2x = 4a \frac{dy}{dx}\] \[\frac{x}{2a} =\frac{dy}{dx}\] Normal is perpendicular to tangent Hence, Slope of Normal is \(-\frac{2a}{x}\) @mimi can you draw the figure? I am confused

OpenStudy (mimi_x3):

lols, im confused as well, i've attempted on drawing it but i think that it's wrong =/

OpenStudy (anonymous):

'From the vertex O a perpendicular OM is drawn to meet the normal at M' Maybe something is not right here, I am not sure but try reading the question again maybe some typo

OpenStudy (anonymous):

ah Okay I think I got it

OpenStudy (anonymous):

maybe the figure is like this